There are six cans of root beer, four cans of ginger ale, and ten cans of cola in a cooler. If two cans are drawn in succession, without replacement, find the probability that the first is root beer and the second is not root beer. Report your answer as a fraction in the lowest terms or as a decimal to the nearest thousandth, Please show me work and formula used thanks so much!
Answer #1

Total beers = 6 + 4 + 10 = 20

Hence,

P(First is root beer and second is not root beer)

= P(1st root beer) * P(Second not root beer | 1st root beer)

= (6/20) * (14/19)

= 21/95